E. M. Facility logo
View this PageEdit this PageUploads to this PageHistory of this PageTop of the SwikiRecent ChangesSearch the SwikiHelp Guide

Updates from the Electron Microscope Facility

REMF home page.
Move information

December 22, 2016.

Tecnai chiller is down. Now it is back!

CPD

November 29, 2016.

The Tecnai is back in service. I have replaced the air compressor with a unit that works well, but is louder than the old one. I will work on getting the old one repaired.
The double tilt holder is in the process of being repaired. I think it may be back this week.
CPD

November 21, 2016.

The Tecnai is down due to a compressor problem. We are working on it. The double tilt analytical sample holder is damaged due to a cross threaded Be hex ring. The Be sample cup has also been twisted requiring replacement of the jewel bearings. It is out for repair.
CPD

November 1, 2016.

New scheduling and billing system is active as of today. The link to Instrument Schedule will now take you to RaDar at https://radar.dartmouth.edu/resources/emlab
To go directly to the scheduling page be sure to log in by clicking on the instrument of interest. After logging in you will be on the instrument scheduling page.
Please let us know if there are problems. COPPS will reman active through November, but billing will be done from RaDar. Please transition to RaDar as soon as possible.
CPD

November 1, 2016.

The Tecnai is back up.
CPD

October 31, 2016.

The Tecnai is down. The stage track ball will only move the sample in the Y direction. It appears the control pad is no good. A new one has been ordered.
The double tilt analytical holder is broken. The Be hex nut has been cross threaded in the Be sample cup. It will have to be sent off for repairs.
CPD

September 22, 2016.

The D3100 and MM8 are back in service. The nanoindenter, too.
CPD

August 29, 2016.

The XL-30 will not pump down. A service call has been placed.
CPD

August 26, 2016.

The XL-30 vacuum is fixed.
CPD

August 19, 2016.

The XL-30 has a vacuum problem. Service is scheduled for Tuesday, August 23.
The Dimension 3100, MultiMode8 and nanoindenter will be out of service for 1-2 weeks due to ventilation work in room 160.
CPD

August 16, 2016.

The XL-30 is back in service, as are all of the other electron microscopes.
CPD

August 15, 2016.

The Horiba Symphony CCD detector is back and in operation. The heat sink fan had failed.
The Tecnai and JEM1010 are both back in service after the weekend power outages. The XL-30 is still baking out prior to restarting the FEG.
CPD

July 29, 2016.

The CCD detector on the Horiba Raman system is not cooling the CCD array. It has been returned for repairs. Only the InGaAs detector is working at the moment.
CPD

July 21, 2016.

The Tecnai is back up. The left hand panel stopped working. Unplugging it and plugging it in fixed the problem.
CPD

July 19, 2016.

The Tecnai is down due to a lens problem. A service call has been initiated.
CPD

July 14, 2016.

The XL-30 low vacuum mode is back in service. One of the two air-activated controllers on the chamber valve failed. It has been replaced.
CPD

June 13, 2016.

The XL-30 will be serviced June 15 and 16. The LV valve will be replaced/rebuilt, and a new scan generator board will be installed.
CPD

June 3, 2016.

The Tecnai is back in service. The XL-30 is also back in service in low vacuum mode.
CPD

May 31, 2016.

The Tecnai is down with an unresolved pole touch. A service call has been initiated.
CPD

May 13, 2016.

The LV mode is not available. A fix is planned forJune 2 and 3.
CPD

May 9, 2016.

The beam blanker and associated electronics have been sent out for testing to see if they are contributing to the e-beam lithography problems. I hope to have them back by the end of the week, or early next week.
CPD

April 1, 2016.

The XL-30 is back in service as of 1pm today.
CPD

March 30, 2016.

The XL-30 is down. The ODP heater is either not heating, or the thermistor is not reading temperature. A service call has been initiated.
CPD

March 11, 2016.

Low vacuum mode is back, and the XL-30 is up and running.
CPD

March 7, 2016.

Chuck will be out of the country from March 14 through March 25. He will have network access during part of this time, and will check email when possible.. Note that there is a +7 hour time difference between Hanover and Kampala, Uganda. For most microscope-related questions, please contact Chris or Louisa.
CPD

March 7, 2016.

The low vacuum mode on the XL-30 is failing. The chamber vacuum is lost when switching to low vacuum. The problem appears to be a bad butterfly valve. A replacement has been ordered, and a service call is being arranged.
CPD.

February 8, 2016.

After applying new thermal paste to the CPU heat sink, the D3100 computer is operational, and the AFM is back in service.
CPD

February 5, 2016.

The Dimension 3100 AFM computer is dead. Do not attempt to turn it on. I will try to revive it Monday morning.
CPD

January 29, 2016.

Tecnai back in service. The XL-30 has a vacuum problem and will be out of service over the weekend and through Monday.
CPD

January 27, 2016.

The 1010 and Tecnai are both back in service. The Tecnai is having routine service Jan. 28 and 29.
The XL-30 will have service Jan. 28.
CPD

January 25, 2016.

The Tecnai is back in service. Remember that it will be down Jan. 28 and 29 for routine service.
CPD

January 21, 2016.

The chiller on the 1010 is down. We hope to have it back in service by the end of the day Monday.
The Tecnai is also down with water flow problems. It will be down until at least Monday evening.
Both TEMs will be off the schedule until they are back up.
CPD

January 19, 2016.

Routine service on JEOL JEM1010 Jan. 19, 20 and 21. If finished sooner, time will be released.
Routine on Tecnai Jan. 28 and 29.
CPD

January 4, 2016.

The FEG tip will be replaced on Jan. 6 and 7, 2016. If all goes well the XL-30 will be back in service by Thursday afternoon.
CPD

December 17, 2015.

The BSE detector is back in service. If this diode is ruined, there will be no more BSE signal.
CPD

December 16, 2016.

The XL-30 is back in service. Please note that a FEG tip exchange will take place in the first week or two of January. This will be a three day service call.
CPD

December 11, 2015.

The XL-30 will be having routine preventive maintenance Dec. 15 and 16. The BSE detector diode is out for repair due to surface scratches on the diode. It should be back in a week or so.
CPD

October 30, 2015.

The Tecnai as back in service. The pole touch problem is more or less fixed. If it comes back, we 'll have to have a more involved service call.
CPD

October 26, 2015.

The CompuStage is giving Pole Touch error, although I don't think there are any. Service is on call to repair. Likely cause is the wires providing 2V signal are loose or broken.
CPD

July 24, 2015.

The Witec confocal Raman microscope is back in service with a refurbished laser.
CPD

June 18, 2015.

The Tecnai is back in service with two new IGPs and a repaired stage.
The Raman laser should be back this week, or next.
Service on the XL-30 to fix vacuum problems June 21 through June 23.
CPD

June 15, 2015.
New IGPs are being installed on the Tecani today through Wednesday. Some stage adjustments will also be made at that time.
CPD

June 1, 2015.

A new objective aperture strip has been put in the XL-30. Currently it is positioned on a 30 micron hole at position 3. Position 4 is also 30microns, 5 is 50 microns, 6 and 7 are both 30 microns.
CPD

June 1, 2015.

The Be hex ring has been found.
CPD

May 29, 1025.

The Be hex ring for the double tilt analytical holder has gone missing. The holder may not be used until a new hex ring is acquired.
CPD

May 29, 2015.

The aperture strip on the XL-30 is dirty. All four 30 micron holes are contaminated. It is now on the 40 micron hole, which will result in higher probe current, but lower ultimate resolution. For most purposes it will not make a difference. For high resolution work and e-beam lithography it will change results. A new aperture strip is on order, and should be in by Monday.
CPD

May 8, 2015.

The Tecnai is back in service as of this afternoon. A new FEG emitter has been installed, the gun electronics have been repaired and the SA aperture drive has been replaced.
New alignments are in the service account for 200kV, 120kV and 100kV. The stage touch alarm has been disabled because it was giving false alarms. Tilt with care. The stage will be pulled and repaired, and IGP3 and IGP4 will be replaced during the week of June 8 (as of now).
CPD

April 30, 2015.

The Raman user is due to ship back on or before June 12.
CPD

April 23, 2015.

The Tecnai will have a FEG tip replacement and other service the week of April 27.
CPD

April 23, 2015.

The Raman laser is out for repair.
CPD

March 31, 2015.

The later on the Raman microscope has stopped working. I am awaiting a quote for refurbishing the unit.
CPD

March 13, 2015.

Power has been restored via a new connection to Remsen. The XL-30 is now available. The Tecnai is still being serviced to correct mechanical alignment problem that was preventing nano probe operation. The JEM1010 is still off from yesterday, and has a problem with the water chiller. A repair order has been started.
CPD

March 12, 2015.

Water leaked into the electrical closet in Gilman. The panel shorted and blew off the doors of both the panel and the door to the room. Our power was thus cut. It has been restored be switching us to a supply from Remsen. All microscopes were down. The Tecnai is still being serviced, so will be down Friday, and probably Monday as well. The XL-30 has just been restarted, but it needs to pump over night at which time we will start the FEG. It is also slated for service through Tuesday to try to address vacuum and beam-realated problems. The JEM1010 is still off. We will try to restart it in the morning. All other instruments are available for use. Sorry for any inconveniences.
CPD

January 21, 2015.

The beam blanker is back on the microscope, and appears to be working properly.
CPD

January 6, 2015.

The beam blanker on the XL-30 has been removed and shipped off for repair. The blanker must have been hit in such a way that it shorts out and is not able to blank the beam. I hope to have it back next week.
CPD

November 18, 2014.

I am off campus. I think the SEM is back up, based on a service report I received. The main valve on the ODP was stuck in the open position. This was repaired and checked, so it should be in service.
CPD

November 14, 2014.

The XL-30 is back in service. The valve was sticking. The set point for the closed position was loose.
louisa

November 10, 2014.

The XL-30 is back in service. The Z-80 processor in the vacuum system had somehow reversed logic so it sensed the main valve was closed when in fact it was open. A new board is on order, but we'll delay installing it until after Thanksgiving as long as operation is normal. Update: The problem is back. I'll reset the board when I get in tomorrow. I have asked that the board be replaced as soon as possible. Update 2: The problem seems to be more than a controller board. The thought now is that the valve is sticking. The engineer will be here Thursday with boards and a solution. Sorry for the ongoing trouble.
CPD

November 6, 2014.

The chiller on the SEM is not working, thus the SEM is not working. Repairs will be made as soon as possible. Update: The pump seems to have stopped. A new one is being sent overnight, so we should be back up sometime tomorrow. yet more updates: The chiller seems to be okay, but still getting VCUDRV down error. Service call has been initiated.
CPD

October 20, 2014.

The address for the FTP server has changed. Now use emlab.dartmouth.edu, rather than remf.dartmouth.edu. All other settings are the same.
CPD

October 6, 2014

We have reset the Windows XP and Windows 2000 computers so they can connect to the E. M. server. Please use the old system of transferring files: winSCP to transfer to remf.dartmouth.edu. Please note that the computers will only connect to remf. They cannot see anything else on the network so you will have to use the remf.dartmouth.edu server for file transfer, not servers you may have set up in your lab. The remf server is visible from on campus. A VPN session is needed for off campus access.
The XL-30 is back to using computer A for saving and exporting images.
CPD

September 29, 2014

The Tecnai is back in service. I have no idea why it turned itself off. FEI thought it might be a bad sensor. We will see.
CPD

September 29, 2014

The Tecnai is back on the fritz. This time it appears to be a bad temperature sensor and switch in the ODP. Service call has been initiated.
CPD

September 25, 2014.

The Tecnai is back in service. The water chiller is repaired, and the ODP heater has been replaced.
CPD

September 12, 2014

The Tecnai is down again. This time it does appear to be an ODP problem. A service call has been initiated.
CPD

September 11, 2014

The Tecnai is back in service. The problem was a bad communication board and a bad connection. Seems to be fine now.
CPD

September11, 2014

The XL-30 is back in operation. The butterfly valve in the ODP was not seating properly. All is well now. Pumping over night to get IGP1 in better state. Good images.
CPD

September 10, 2014

The XL-30 vacuum problem is back. A service call has been initiated.
CPD

September 9, 2014

The XL-30 is back in service. The vacuum manifold was leaking. After rebuilding the maniforls all is well.
CPD

September 5, 2014

The XL-30 vacuum system is down. Service is set for Monday, Sept. 8.
CPD

August 27, 2014

A power outage is tentatively planned for Sept. 3 from 9:30 am until 11:30 am. The microscopes will be turned off at 9 am. With any luck we will be back in business by mid-afternoon. However, I have blocked out the whole day to ensure that all is well.
CPD

August 27, 2014

The Tecnai has a bigger problem than the ODP. The entire vacuum system has stopped responding to the computer/electronics. It is being worked on now. Time to being back in service is uncertain. I will update as I have news.
CPD

August 25, 2014

The ODP on the Tecnai has a burned out heater. Service call set. Will be back on line as soon as heater replaced.
CPD

August 8, 2014

XL30 is back on-line

August 7, 2014

The xl30 is undergoing service tomorrow. New diffusion pump to fix vacuum problem.


August 6, 2014.

The XL-30 has a new objective aperture strip installed. The lower column was purged with Nitrogen gas for 15 minutes. The beam blanker was cleaned, as was the pole piece insert. The banker was re-aligned. The microscope has been aligned at 20kV and 10kV, so should be in excellent shape for all uses.
A replacement of the ODP and associated valves (the parts causing the vacuum error on vent) is in the works, but the date is uncertain at this point. It should involve no more than one day down time.
CPD

July 31, 2014

File transfer instructions for various machines in the E. M. Facility

1. Image and other data from the XL-30.

In the near term we will have to use a USB drive (the red one that should be left under the XL-30 monitor) to transfer images from the XL-30 computer, the HKL computer (A) and the EDAX (D)computer to the NPGS computer (C). Do not use any other USB drive.

Images should be saved to the XL-30 computer in the C:\\Archive\Fei\Fei.mdb database (it should be open unless there has been a reboot of the XL-30 computer). Export images from the database to this same folder, then copy them to the USB drive. Then move the images to the NPGS computer (C). Then use winSCP as usual to move them to the server. If there is not enough space on the USB drive, delete some of the files and folders.

Eventually we will have three computers connected to the NPGS computer via ethernet and will not need to use the USB drive.

2. Images from the JEM1010 TEM.

There are two computers in the room: The AMT camera control computer (Windows XP)and the Windows 7 computer (W7). The XP computer is not connected to the internet. It is connected by a support network to the W7 machine, which is connected to the internet.

There is a KVM switch that allows you to toggle from one machine to the other using the same keyboard, video and mouse. The switch is the small black button to the left of the monitor stand.

The W7 machine should always be on. The AMT machine needs to be started at each microscopy session. Start and log in as EM Lab.
Use the AMT machine and image acquisition as usual. When you are finished, close the AMT program and then copy files from the XP machine to the red USB drive. Remove the USB drive from the AMT computer (XP) and insert into the W7 computer. Toggle to the W7 machine and use winSCP to transfer files to the server. Turn off the AMT computer, leaving the W7 machine on. Remove the USB drive and leave on the keyboard.

3. File transfer from the Ubi-1 nanoindenter or Dimension 3100 AFM

The Ubi-1 and D3100 have Windows XP computers that have to be off the network. They are now connected through a bridge computer running Windows 7 which is on the network. Data from the Ubi-1 and the D3100 need to be saved to a shared folder on the Windows 7 machine. The following are the procedures for connecting and transferring data.

We now use the left hand monitor and the mouse and keyboard to connect to all three computers. To move from one to the next press in sequence control, then alt, then shift and 1, 3 or 4.
The bridge computer is 3, the Ubi-1 computer is 4, the D3100 computer is 1 and there is no computer on 2.

1. Be sure the bridge computer is on (it is labeled number 3).
2. Start the computer you will use (Ubi-1 or D3100). Log in as usual.
3. When the computer is fully started, open the desktop shortcut to the shared public folder. You may be warned that not all network drives could be mounted.
4. When prompted for a password for the shared drive, enter the same password as the one used for file transfer. The public folder should open.
5. When acquiring data, be sure to save them to the shared folder on the bridge computer.
6. When finished, toggle to the bridge computer (control-alt-shift-3) and use winsSCP to move file to the server.
7. Leave the bridge computer on. Turn off the other computer as usual


4. File transfer from the Witec CRM200 confocal Raman light microscope.

The Witec machine (XP) is connected to a Windows 7 machine (W7). To toggle from one to the other press the Scroll Lock key twice. The keyboard has a KVM mark above the key to make it easy to find. The W7 machine should always be on. Turn on the XP machine and before starting the Witec software be sure to open the link too open the shared folder on the W7 machine the desktop. From the Witec software you should be able to save to this folder. When finished, toggle to the W7 machine and use winSCP to move files to the server.

CPD

July 29, 2014

The Tecnai has a new FEG. Gun settings are in the service account, or after tomorrow, in my account. Other alignments should not be changed.
CPD

July 29, 2014

The venting problem has been resolved. The leaking pressure gauge on the air line lead to oil contamination of the ODP butterfly valve. This has been cleaned up, and the venting works well. A new ODP may be installed on 7-30-14, if one is in stock.
CPD

July 25, 2014

Some of you know that we have been having trouble with the SEM vacuum system. When you click the Vent button the VCU driver stops so you cannot vent. Sometimes we also get a odp critical back pressure error, too. In either case, you can proceed by clearing the errors, then pressing the VAC button on the microscope - one of the 4 lighted buttons. You may have to clear and press the VAC button as many as ten times to get it back in operation. Once this is done everything seems to be fine. You will also have to press the HT button when the vacuum is back on.
A service call is scheduled for Monday.

CPD

July10,2014

JEOL TEM 1010 Users: The TEM 1010 specimen holder is being stored in desiccator (Large, tan JEOL EMDSC-U10A) located at TEM 1010 entrance doorway.

To use:
1. make sure POWER is on, If not Press on POWER button
2. Press VENT button (@1-2 minutes to reach atmosphere).
3. open door, remove holder.
4. Close door; push PUMP button on. Takes several minutes for vacuum to reach into the green. If it doesn’t reach into the green; VENT, check O-ring on door for dust. Then, while pressing lightly on the door to make sure it seals, PUMP down again.
5. At end of session store Specimen holder back in the desiccator.


July 3, 2014.

The Tecnai needs a new FEG. This will take place in the next week or two. The actual exchange will take the better part of a week to complete. I will post a schedule when I have it.
CPD.

June 2, 2014.

The stray specimen was removed, but the objective aperture strip was ruined in the process. A new strip will be in Tuesday. I hope a service engineer will also show up.
CPD

May 28, 2014.

The Technai is down due to an errant specimen. We will need a service call to extract the sample and reset the stage. Scheduling for service is in progress.
CPD

April 24, 2014.

File transfer from the Ubi-1 or D3100 AFM.

The Ubi-1 and D3100 have Windows XP computers that have to be off the network. They are now connected through a bridge computer running Windows 7 which is on the network. Data from the Ubi-1 and the D3100 need to be saved to a shared folder on the Windows 7 machine. The following are the procedures for connecting and transferring data.

We now use the left hand monitor and the mouse and keyboard to connect to all three computers. To move from one to the next press in sequence control, then alt, then shift and 1, 3 or 4.
The bridge computer is 3, the Ubi-1 computer is 4, the D3100 computer is 1 and there is no computer on 2.

1. Be sure the bridge computer is on (it is labeled number 3).
2. Start the computer you will use (Ubi-1 or D3100). Log in as usual.
3. When the computer is fully started, open the desktop shortcut to the shared public folder. You may be warned that not all network drives could be mounted.
4. When prompted for a password for the shared drive, enter the same password as the one used for file transfer. The public folder should open.
5. When acquiring data, be sure to save them to the shared folder on the bridge computer.
6. When finished, toggle to the bridge computer (control-alt-shift-3) and use winsSCP to move file to the server.
7. Leave the bridge computer on. Turn off the other computer as usual

CPD


January 23, 2014.

A new aperture strip has been installed in the XL-30. The lower chamber has also been purged with dry Nitrogen. The gun was adjusted at 20kV and 10kV. The old aperture strip was contaminated with particulates (see picture below). Anyone examining small particles of any kind must ensure they are well attached to the mount. All particulates must be coated. Finally, the low vacuum insert must be used when examining particulates even if you only use high vacuum mode.
CPD
Uploaded Image: aperture-3.jpg

January 7, 2014.

The power outage killed the HD on the Tecnai. A backup disk is in use, but without the latest alignments. I will redo the 200kV and 80kV alignments as soon as possible.
CPD

January 6, 2014.

The recent power outage has killed the stage power supply. As a result the XL-30 is down until tomorrow afternoon. I will update as soon as the power supply is replaced.
CPD

December 23,2013.

Readjusted apertures and checked alignments @20kV. Good images. Beam current 14.59 pA @ 20kV, spot 1.
CPD

December 9, 2013.

There have been some imaging problems with the XL-30. I have contacted the service engineer, and have made some evaluations. This morning the microscope is working perfectly. I have, however, found several stage rotation positions where the ground is not good. I will have these addressed in a service call. I also found a piece of the backing paper from the mounting tape in the microscope. The beam blanker seems to be working well, and does not appear to have any effect on the image. At this point all I can say is that the machine looks as if it is working well. Make sure the stage is at zero rotation. You should probably home the stage when you load the samples to make sure it is at zero where there is a good ground.
CPD

December 3, 2013.

The XL-30 has had the lower column liner tube and surrounding areas cleaned. The beam blanker was also cleaned. Images are good at 80,000x at 15kV, spots 1 and 3.
If you are looking at small particulates, you must use the low vacuum column insert, even if you use high vacuum conditions.
CPD

October 28, 2013.

The XL-30 service is continuing. Today we are replacing the feed through fro the large field detector. It has been arcing, which causes streaks in the GSE images. A new aperture strip was installed and the beam blanker adjusted. Should be back in service by noon.
CPD

October 24, 2013.

The XL-30 is not usable. It appears that there is contamination in the column, the scintillator needs replacing and there have been some computer issues. A service call is scheduled for Friday, Oct. 25.
CPD

July 25, 2013.

The SEM is back in service with a new valve block and a new FEG. The cold stage will be installed as soon as I get the tubing for the gas connections.
CPD

July 23, 2013.

We are almost out of the woods with the XL-30. Vacuum is back, charging is gone, but now the FEG is acting up. A new FEG will be here tomorrow. If we are lucky, the current one will recover over night and the microscope will be ready at the end of the day. If it is not better (restored emission), then the new one will have to be put in with an over night bake out. So, no SEM until Wednesday evening or Thursday. Worst case is it will be back only on Friday.
I will delete sessions for tomorrow. They will have to be rescheduled. I'll leave the Thursday sessions active until I know when we will be back on line.
I am sorry for this avalanche of bad luck and delay, but the end is near. For those of you interested, the cold stage is back, as is the heating stage. I will install and test both next week.
CPD

July 19, 1023.

The XL-30 SEM is still down. The valve block for the column isolation valve was successfully replaced, but there is some kind of charging in the column, so it is not possible to use the SEM. In trying to resolve the charging problem the microscope has been opened again, but now they are having trouble getting it to pump. All of this means the XL-30 will be down at least through Monday.
Late update: The microscope will be down until at least Tuesday.
CPD

July 16, 2013.

The focus knob has been replaced. It may feel a bit odd since it is from a slightly different model microscope. It should work.
CPD

July 15, 2013.

The fine focus knob on the 1010 is not working, thus the microscope is not working. Repair call has been made. JEOL are waiting for a new knob.
CPD

July 10, 2013.

The XL-30 service has been moved to July 16 and 17.
CPD

July 8, 2013.

The XL-30 column valve service will be Thursday and Friday of this week. The microscope should be back in operation by Friday at 4:30pm or before.
CPD

July 8, 2013.

The NPGS system should be in working condition at this point. Note that the new computer is Windows 7. Please do not make any changes to the computer. The XL-30 is having a column valve issue, so we will be having it serviced this week (I don't know the date yet). The servicing will take two days - one for the repair the, second for bake out and restart.
CPD

June 20, 2013.

There will be scheduled service on the XL-30 on June 27 and 28. The column valve seems to be leaking slightly, which degrades the vacuum. This is a quick fix, but requires a bake out of the IGPs, so it will take the full two days.
CPD

June 19, 2013.

The Tecnai is back in service. Users will have to update alignments either on their own, or by borrowing recent sets from me.
The XL-30 is also back in service.
The NPGS computer is dead. A new computer has been ordered. As soon as it arrives and is set up I will let users know.
CPD

June 18, 2013.

The Tecnai was coming along nicely until the morning power outage. It has recovered, but now has had two discharge events in the gun area. It may not be ready for another day or two.
CPD

June 18, 2013.

The XL-30 is back up. It should be ready to go at 1pm. There may be some problems with the XL-Docu database, but I will try to resolve those soon. I had to rebuild the computer drive from the Ghost image of the drive.
The NPGS computer appears to be completely dead. I will try to get a new machine as soon as possible.
CPD

June 18, 2013.

There was a power outage early this morning in the EM wing of Remsen. The computer controlling the SEM is down as is the NPGS computer. I will get things fixed as soon as possible.
CPD

June 14, 2013.

The Tecnai now has a new FEG source. However, the blocked beam was still evident. A tiny piece of metal was found in the column liner near the C1 fixed aperture. This has been removed, and the TEM is pumping over the weekend. We will know Monday afternoon if that has resolved the problem.
CPD

June 3, 2013.

The Tecnai is again down. There is something blocking the beam high in the column. It may be a tungsten oxide from the FEG suppressor. We should know shortly how we will proceed to get the Tecnai back in service.
CPD

May 31, 2013.

The Tecnai is back in service. You may need to adjust your low magnification image alignments, since the lower column was in pieces during service.
CPD

May 24, 2013.

The Tecnai service will extend into the middle of next week - to May 29 or so.
CPD

May 20, 2013.

The Tecnai will be undergoing service for vacuum and lower column alignment on Tuesday and Wednesday of this week. It should be back in service by the end of the day Wednesday.
CPD

May 3, 2013.

XL-30 back in service. Vacuum loss due to leaking cold stage line during use resulted in loss of oil from ODP. That in turn prevented the ODP to heat properly. Oil level adjusted, and back in business.
CPD

May 2, 2013.

The XL-30 is down. Vacuum problem. Under investigation.
CPD

March 25, 2013.

The Tecnai is back up. The lower camera should work. If there are problems, let us know. The airlock has also been replaced, so there should not be vacuum dumps during specimen changes.
CPD

March 14, 2013.

The Tecnai is down due to a failed air compressor. A new unit will be acquired as soon as possible.
CPD

March 12, 2013.

The 1010 TEM is down. The beam has vanished. We will try to get it back in service soon.
CPD

March6, 2013.

The Tecnai is back in service.
CPD

March 5, 2013.

The Tecnai is still down. The stage insertion tube has been replaced. A couple of vacuum problems have been solved. Now the stage is being re-calibrated. This will take until the end of the day Wednesday.
CPD

March 1, 2013.

The Tecnai air lock needs a repair. One of the inner mechanisms needs to be replaced. Repairs should be complete on Monday.
CPD

January 9, 2013.

The digital cameras on the Tecnai are down while we figure out what is wrong with the computer. Updates to follow as needed.
CPD.

December 20, 2012.

The E. M. Facility will be closed from 12-22-12 through 01-01-13, college winter break. During that time there will be no liquid nitrogen delivery. The X-ray detectors will be out of service during their annual defrosting cycle.
CPD

December 13, 2012.

Update (100am): The Tecnai is now pumping. It should be ready to go this afternoon.
The vacuum on the Tecnai was dumped by a user on Thursday. The microscope will not pump down properly. A service call has been started.
CPD

December 7, 2012.

The Tecnai is back in service. A set of alignments (CPD-B12-12) has been saved in Daghlian folder if you have trouble with yours. The AMT software is again not reading diffraction camera lengths. I will get that fixed next week.
CPD

December 4, 2012.

A loaner bottom camera is on order for the Tecnai. It should arrive in a day or two. A new screen for the side-mount camera is also on order. Our bottom camera is on its way to AMT for evaluation and repair.
CPD

December 3, 2012.

The Two cameras on the Tecnai are out of service. The screen on the side-mount was scratched during service. We are waiting for a new screen. The bottom-mount camera is malfunctioning - vertical lines, no usable image. A service call is in progress. The Tecnai itself seems to be okay.
CPD

November 30, 2012.

The Tecnai should be back in service 12-03-12 at noon. I will reset the basic 100kV and 200kV alignments. The screen on the side mount camera is scratched, so only the bottom mount is available at the moment.
The x-axis on the XL-30 stage drive has a loose coupling, so it may throw a Tracking Error. It can be used manually, so work around it. Service is expected on Tuesday.
CPD

November 27, 2012.

The Tecnai is still down. Lenses have been replaced, the objective stigmator has been repaired. Now we are trying to get the upper gun tilt coil to move from -13V. With luck it will be back in service by the end of the week.
CPD

November 16, 2012.

The Tecnai is almost ready to back into service. There is a problem with the objective stigmators. They don't seem to do much. We do have new P1 and P2 lens coils, and a new lens power supply set.
The XL-30 is back in service.
CPD

November 7, 2012.

The Tecnai continues to have problems. The damaged lenses were replaced and the system has pumped down, however
there is a small leak in the camera/viewing chamber that causes the buffer cycle to turn on every few minutes. To add to the problems, one of the replaced
lens modules is no good and needs to be replaced. It is unlikely that it will be back on line before Tuesday or Wednesday of next week.
Chuck here for updates.
CPD

November 5, 2012.

The Tecnai is down. The lens cooling is fixed, and while the two damaged lenses have been replaced, there is now a vacuum problem.
That should be sorted out today. Then we need to bring the HT back and get the FEG operating. With luck this can all take place no later than Wednesday.
CPD

October 31, 2012.

The Tecnai has lost several lenses, among other problems. It will be down until at least the end of this week (11-02-12).
CPD

October 30, 2012.

Routine service for the XL-30 has been scheduled for November 7 and 8. If all goes well the instrument will be available for use at 5pm on Nov. 8.
CPD

October 26, 2012.

Routine service is planned for the XL-30 during the first week in November. Details as they develop.
A service call has been initiated for the Tecnai. There may be lens control board issues.
CPD

June 21, 2012.

All the instruments are back on line. The Tecnai will be shut off Monday, June 25, for much of the day as we trouble shoot the UPS unit that has failed.
CPD

June 20, 2012.

We have had yet another power outage. The 1010 is being restarted, as is the XL-30. The XL-30 may need a bake out. We will know by 10am. The Tecnai is down until the faulty UPS unit can be repaired or bypassed.
Update: The 1010 appears to be starting easily. The Tecnai is also restarting. It will be shut down Monday, June 25, for a few hours so we can have the UPS serviced. It will take most of the rest of Monday to restart it. The XL-30 is in bake out mode. I will try to restart the FEG around 3pm.
Update 2: The Tecnai is back up, and will be ready to use in the morning (6-21-12). The XL-30 has finished a 4 hour bake out. After cooling, we will restart the FEG. It will also be available for use in the morning (6-21-12).
CPD

June 18, 2012.

The Tecnai is down due to a failed computer. Service call is being scheduled.
CPD

June 11, 2012.

There was a short power outage this morning. All microscopes are back up.
CPD

June 7, 2012.

The XL-30 cold stage has been repaired. Users are being too rough on this device. The threads to the hose barbs were crushed. If the nuts do not go on smoothly stop and contact me. The tips of the hose barbs on the cold stage are partially crushed as if they have been dropped. Please be careful. A replacement stage will be very expensive.
CPD

June 6, 2012.

The SEM cold stage is out of service. The threads on the hose barbs have been damaged. I will try to have them repaired in the next day or two.
CPD

June 5, 2012.

Both the Tecnai and XL-30 are bock in service.
CPD

June 4, 2012.

The XL-30 is baking and should be ready tomorrow mid-day. The Tecnai is back up, but needs to pump until Tuesday.
CPD

June 4, 2012.

After the two power outages on Saturday the Tecnai and XL-30 are both down. There appears to be a leak or other problem in the specimen airlock on the Tecnai such that it cannot pump the column. A service call is in, and we will have someone on site by Wednesday, if not sooner.
The XL-30 was off long enough to require a bake out of the IGPs overnight. With luck it will be back in service by tomorrow mid-day.
CPD

June 1, 2012.

The XL-30 SEM will have two days of preventive maintenance on June 7 and 8. It is possible it may not be available over the June 9-10 weekend. Check here and on the schedule for updates.
CPD

May 16, 2012.

JEOL 1010 and Tecnai back in service. The 1010 will get a new video card tomorrow morning. The XL-30 has been baked out and the FEG restarted. In the morning it will be put back in service.
CPD


May 16, 2012.

JEOL 1010 back on line, but over night power outage has all microscopes down for at least morning of May 16.
CPD

May 1, 2012.

The JEOL 1010 TEM is down. The specimen holder was incorrectly inserted resulting a vacuum loss and catastrophic shut down. We are not able to get the machine going. JEOL technical service has been contacted to try to determine the nature of the problem.
CPD

February 7, 2012.

XL-30 service did not resolve vacuum problem. It appears that there is either a fault in one of the vacuum control boards, or a sensor problem. Adjusting the air pressure may have helped. Further investigation is in progress. In the short term we will have to manually clear the errors and reset the vacuum if the error on vent occurs.
CPD

February 2, 2012.

Service on the XL-30 on Feb. 6 to resolve a recurring vacuum error.
CPD

Novermber 30, 2011.

The EDAX X-ray detector is back in service.
CPD

November 10, 2011.

The EBSD system is back in service. The new screen has already been damaged. I will replace it if it interferes with operations.
The X-ray detector has been repaired and is in line for testing and should be back in a week or two.
CPD

October 24, 2011.

The SEM will be out of service until Wednesday, Oct. 26. A new FEG emitter is being installed, and a chamber vacuum issue is being addressed.
CPD

October 5, 2011.

The XL-30 now appears to be okay. Yesterday the upper IGP showed a bad vacuum. Today it is fine and the microscope seems to work well.
CPD.

October 4, 2011.

The XL-30 SEM is down due to an excessive pressure in the upper IGP. A bake out will be done on 10-05-11.
CPD.

September 29, 2011.

The screen on the Nordlys EBSD detector has been scratched and needs to be replaced. Parts have been ordered, but will take two to three weeks to arrive.
CPD.

August 30, 2011.

The microscopes are all back up. If you need to use the XL-30 at less than 5kv, let me know. We will need to replace the FEG soon, but I would like to delay another month or so, if possible.
CPD.

August 29, 2011.

Power outage over the weekend due to Irene. The Tecnai is back up. The 1010 is on, but wont be available until late Tuesday. The XL-30 is in bake out mod an will, with luck, be ready by noon Tuesday.
CPD.


July 29, 2011.

For the second day in a row there has been a power outage. The Tecnai is back in service. The 1010 is on the way back up, but will probably not be ready until later this afternoon. The XL-30 needs to get the lower IGP to <2x10-7 mBar. If it is not there by noon, I will have to do a short bake out.
CPD

July 20, 2011.

The XL-30 is still down. The problem is getting the IGPs to restart. I think they may be in good shape now, but will not know until the end of the day.
CPD

July 19, 2011.

The bake out failed, so the XL-30 will be out of service until Wednesday at noon at the earliest.
CPD.

July 15, 2011.

The XL-30 requires a bake out over the weekend and Monday. If all goes well it will be back up on Tuesday at noon.
CPD

July 13, 2011.

The power outage has stopped the XL-30. It may not be back on line until late Friday. Other EMs are backup, or on the way back.
CPD

June 17, 2011.

The JEOL 1010 TEM has a loaner goniometer while ours is being repaired. Check with Louisa regarding status.
The XL-30 SEM has a cracked BSE detector diode. The A half is still operational. On flat samples it should be fine. A new diode will be ordered in July.
CPD

May 23, 2011.

Planned service for both the XL-30 and Tecnai. XL-30 service planned for June 9 and 10.
Tecnai routine service planned for May 26 and 27.
CPD

May 23, 2011.

The XL-30 is back in service. There was a bad start capacitor.
The Tecnai has a stuck airlock. It is not possible to put a holder into the column. Service is scheduled for Tuesday.
CPD

May 19, 2011.

The SEM is down due to a broken water chiller. I hope to have it back in service sometime tomorrow.
CPD

May 10, 2011.

The SEM is back in operation - has been since last week. Routine service due in early June.
CPD

May 2, 2011.

The column isolation valve was dirty. It has been cleaned. Now we are baking the column. Tuesday morning I will see if the vacuum holds on venting. If so, we are back in business. If not, we'll have to open the column again and clean out the lower section.
CPD

April 29, 2011.

After several bakes outs of the IGPs, the status of the microscope is that the FEG is OK. There appears to be a leak in the column isolation valve. On venting the lower IGP pressure goes from ~3.5x10-8mBar to ~1.7x10-7mBar. When pumping down, the vacuum returns to the 10-8mBar range. Monday we will have service to clean the isolation valve. At that time a new scintillator will also be installed. This should improve secondary electron signal strength. With luck, we'll be back in operation by the end of the day Monday.
CPD

April 26, 2011.

The XL-30 SEM is down. There is a higher than normal vacuum reading in the upper IGP. This is either due to contamination in the coulomb tube that is outgassing, or a damaged FEG. In either case we need a service call. As I know more, I will post new updates.
CPD

February 16, 2011.

The XL-30 SEM will have a service call to fix the LV mode pump failure. This should be a one day fix. Other modes (HV) are fine.
CPD

January 24, 2011.

The BSE detector is back in service.
CPD

January 21, 2011.

The BSE detector on the XL-30 is broken. It appears to be wires that have been pulled loose. It should be repaired Monday, Jan. 24. Please use care when installing and removing.
CPD

January 7, 2011.

There will be yet another planned power outage on Monday, Jan. 10 from 7:30 to 8:00 am. The microscopes will take some time to get back up. They have been reserved until noon. Check the schedule to see if they are back sooner.
CPD.

November 15, 2010.

There will be a power outage Tuesday, 11-16-2010 starting at 7am. The E. M. Facility will shut down the machines by 5pm on 11-15-2010. Instruments should be back up by late afternoon on 11-16, if all goes well.
CPD.

November 8, 2010.

A new screen has been installed in the EBSP detector. A cover is being made that needs to be put over the port to the EBSP detector when it is not in use.
CPD

November 8, 2010.

XL-30 SEM is back to having a 30 micron aperture as the default.
CPD

October 20, 2010.

The FEG on the Tecnai will be replaced the week of November 1. The microscope should be down for the entire week. The beam is somewhat unstable at the moment, so critical work should be done after the new FEG is installed.
CPD

October 18, 2010.

A new 40 micron aperture has been placed in the XL-30 SEM. For most users this should be the aperture of choice. For lithography, it will roughly double the probe current of the 30 micron aperture. If you experience any problems, please let me know. If you need to do very high resolution imaging, or are working with beam sensitive specimens, you may want to use the 30 micron aperture. Probe currents for both the 30 and 40 micron apertures at 20KV are listed here http://remf.dartmouth.edu:8080/NPGS/2.
CPD

October 18, 2010.

The NSIV controller has been sent back to Veeco (now Bruker) for repairs.
CPD

September 2, 2010.

We have ordered and received the parts to be replaced in the NanoScope IV controller. The repair will take about a week, since there are surface mount components that have to be replaced. With luck we will be back in business in about 7-10 days.
CPD

August 20, 2010.

The current thought on the AFM trouble is that the DSP chip, or an associated op-amp, is burned out. I will order new parts when I get the parts list. Then we'll see if I can swap out the correct part. With a bit of luck it will be operational by the end of next week.
CPD

August 17, 2010.

I am trying to figure out the current problem with the AFM. So far I have received and replaced the back board of the 3100 with no change in operation. I will try plugging in the MultiMode to see if the controller is even reading the signals from the AFM head. If it is, then I will have to return the loaner board, our board and the Dimension AFM head for testing and repair. If the controller is not reading thengs it will either have to go in for repairs, or I will have to try swapping out boards from Veeco. I will keep you posted.
CPD

August 12, 2010.

More problems with the AFM. Now it appears there is a failed board in the Nanoscope controller. A loaner board is on its way. With luck that will fix things and we will be back in service.
CPD

August 10, 2010.

The double tilt holder has to be sent off for repairs.
CPD

July 27, 2010.

The double tilt holder for the Tecnai is generating a tracking error. FEI service thinks it may be a stuck potentiometer in the holder, but it is not something I can fix. We will have a service call to determine the actual problem. I will keep you posted.
CPD

July 15, 2010.

At last the AFM is back in service. The new computer is set to go. Old data files and workspaces have been restored. Please see me prior to use.
CPD

July 1, 2010.

Revised rates are in effect as of today. See the Fees page on this site or for a full listing see the E. M. Facility Service Catalogue page in the NCCC Core Operations web site.
XL-30 internal rate is now $40 per hour. The outside rate is $280 per hour with operator.
TEM time is now $45 per hour internal rate, $280 per hour outside rate.
technician time is now $48 per hour internal rate, $150 per hour external rate.
CPD

June 7, 2010.

A replacement computer for the AFM is scheduled to ship from Santa Barbara on June 17. As soon as it arrives I will set it up.
CPD


June 4, 2010.

The BSE detector is back in service. Please be careful installing and removing the detector. The spring clips must not be unsprung!
CPD

May 17,2010.

A replacement diode for the BSE detector is scheduled to be delivered by late June.
CPD

May 10, 2010.

The BSE detector on the XL-30 is broken. Until I can find a replacement, we will be without this imaging mode.
CPD

May 5, 2010.

The XL-30 is back in service. We have made a couple of changes to the system. The XL-30 computer is no longer connected to the network. It is directly connected to the HKl computer (switch A on the KVM switch). There is a shared folder there into which all image data must be saved. When you export images from the database you need to create a folder on the shared volume. The HKL computer is connected to the network through a second ethernet card. Move files with WinSCP as before. If you have trouble, let me know.
CPD

May 4, 2010.

The XL-30, which has been having some computer issues (magenta screen, freezes, etc.) has now given up the ghost. It may just be a bad power supply. We'll try to get things back on line as soon as possible. The service engineer will be here tomorrow morning. The good news is that it is clearly just the computer, not the microscope. Check here for updates.
CPD

April 28, 2010.

There is continuing trouble with the AFM. The latest view from Veeco is that the comnputer needs new memory. I am not sure this is the problem, but I will replace the memory and see how it goes. If this does not fix the problem it will cost about $5,500 to get a new computer from Veeco set up to operate the controller. That is a lot for a computer, but it is because they have a supply of the now antiquated systems needed to use their boards and software. I have looked into getting a cmputer directly, but it is hard to find the correct configuration (and Veeco is not great about giving specifications - the ISA bus needed for their custom DSP board is one problem). An additional problem is that I have no funds to replace the computer. If it need to be replaced, I will come to you hat in hand. A final problem is that I think there may be a problem with the scanner itself, although I cannot be sure of this until the computer is back in service.
CPD

April 15, 2010.

The AFM is back in service, but with some changes in software.The AFM is back in service, but with a few changes. For some reason the version
6 software will not run. You will need to use version 5. There is a shortcut to
it on the desktop calles SPM. All of the functions are still present, but in
different places. Please see me for instructions, if you need them. There is now
one capture folder on the D drive. After capturing, either move your images to a
separate folder of your own, or delete after transferring.

To tranfer files you must use the psftp.exe program from its short cut on the
desktop.

Do the following:
1. type "open remf.dartmouth.edu"
2. log in as em lab with password micronx
3. Go to your folder on the server by typing "cd /Volumes/Mongo/public/your
folder/"
4. Move files by typing"mput D:\capture\your folder\*.*" If you named the files
with a root name you could type "rootname.*" to transfer just those files. The
files will be listed as they are moved.
5. Type "exit"

Note that BlitzMail is not on the computer. Please do not place a copy there.
CPD

April 5, 2010.

A new copy of frame grabber board drivers is on its way. That should fix our problems and give us a backup of the drivers.
CPD

March 30, 2010.

The AFM is down due to a problem initializing the stage. This is a software problem. We are working with Veeco to resolve the trouble.
The XL-30 will have routine service on April 8 and 9. Please plan accordingly.
CPD

March 22, 2010.

XL-30 service cancelled. Will be rescheduled in the next few weeks.
CPD

March 2, 2010.

The JEM 1010, the XL-30 and the JSM5310 will all have routine service during the inter-session period March . Please consult the schedules for details. We have reserved three or more days for the service. If service is finished early, time will be released. Please plan ahead.
CPD

December 22, 2009.

The E. M. facility will be closed during the college winter break, 12-24-09 through 01-03-10. We will open again on Monday, Jan. 4, 2010. Please note that even if you come in to use the instruments, there will be no liquid Nitrogen for cold traps. The X-ray detectors will be at room temperature during the break in our annual drying out period. This will prevent water and ice accumulation in the detectors.
CPD

November 24, 2009.

The JEOL 1010 is back in service.
CPD

November 13, 2009.

New pump and valve have arrived and been installed. The water is flowing, but the TEM is still down. We have a service call to reset the HV and other boards in the microscope. The 1010 will probably not be up before Tuesday of Wednesday.
CPD

November 12, 2009.

The JEOL JEM 1010 TEM is down due to a failed water pump. Parts have been ordered and should arrive Friday. With luck we will have it back on line by Monday afternoon.
CPD

November 2, 2009.

The Tecnai has a water leak in the manifold stand. This causes the chiller to lose water and eventually shut down, not to mention is inherently unsafe. The repair Tuesday should be quick. A new fitting needs to be installed to replace the leaking one.
CPD

September 28, 2009.

The wires to the heater and thermocouple on the Gatan C1002 cold stage have broken. Both items need to be replaced. An order is in, but will take several weeks to be filled.
CPD

September 8, 2009.

The XL-30 FEG tip change went well. Imaging is excellent. Only current problem is getting rid of a virus from the control computer. For image transfer it is essential that only SFTP or burning a CD be used. Do not mount remote disks or file sharing folders.
CPD

August 27, 2009.

The XL-30 SEM will have the FEG tip changed on September 3 and 4. This will require an overnight bakeout and re-alignment the next day. If all goes well the SEM should be back inservice by the evening of 9-4-09.
CPD

August 11, 2009.

All instruments are back in working order.
CPD

August 10, 2009.

XL-30 computer is down. Hard drive is dead. A new unit is on order and should arrive tomorrow morning. Check back for updates.
CPD

August 7, 2009.

The XL-30 high definition image acquisition is back in service. The cable to the frame grabber was not mating properly.
The cold stage heater wires have broken and need to be repaired. Out of service until next week.
CPD

August 5, 2009.

The new cable for the HD images did not fix the problem. A service call has been scheduled for Friday, August 6.
There was a blue screen of death on the HKL computer this morning. After several restarts in safe mode and a full diagnostic test, it appears to be working. If the blue screen occurs again the unit will go in for repairs.
CPD

July 31, 2009.

XL-30 HD image capture does not work. New cable on order.
CPD

March 18, 2009.

The Dimension 3100 AFM is down for a piezo tube repair. No sign ups will be allowed until it is back in service. The estimate for repair and return is two weeks from the time Veeco receives the unit. It may be less than this.
CPD

December 19, 2008.

A new series of snowflake images have been placed on the page.
The BSE detector has been replaced, and the EDS system is back in service.
The lab will be closed during winter break.
CPD

March 13, 2008.

EMFUG Luncheon is scheduled for Tuesday, April 15 at the Inn.

The X-ray detector on the XL-30 has a pinhole leak and has been removed for repairs. I am not sure how long it will be gone, but will update you when I know.

We are also still trying to resolve vibration problems with the microscope. You may not have noticed the problem if you have not tried to do high resolution imaging. There are a couple of potential solutions. We may need a new stage, but it may also just be a problem with some components. If mechanical problems with the stage are ruled out, we'll have to have a new site survey to see if new fields or vibrations have developed since installation of the microscope.

For lithography, imaging and EBSD there should be no problems.
CPD

February 28, 2008.

The XL-30 appears to have blown another aperture. We are trying to clean some old aperture strips, and are arranging to either obtain a disk aperture holder or machine one. We will also be getting a new control computer for the XL-30 to resolve some of the freezes and magenta screen issues.
CPD

February 11, 2008.

We are in the middle of a FEG change today through Wednesday noon. Also being addressed are problems with the stage and an overheating computer (due to fan problems).
CPD

July 25, 2007.

This from Facilities:
On August 2, 2007 Dartmouth College will start the construction of a underground utility corridor on the North Campus.

During the week of August 6 and 13, the Contractor will start the installation
of three (3) steam utility manholes located adjacent Kellogg Hall, at the
northeast corner of the Building 37 parking lot and near the entrance into the
Dewey Field Parking Lot.

This will certainly cause some vibrations.

December 6, 2006.

This from DMS Facilities regarding the pile driving:
From: Jon Rader
DMS - Facilities

Date: December 6, 2006

Subj: Construction Project - Level 1 Vail

I have been in contact with the general contractor for the Vail Electrical project trying to work blocks of time when the pile driving work could be shutdown so you could use your scopes. The best that I can do is to have the driving work shut down at 3:30 every afternoon. This schedule will go into effect on Monday, 12/11. The pile driving is suppose to be completed by Friday, 12/15.

Jon Rader
DMS - Facilities

November 20, 2006.

The following note from DMS Facilities outlines some construction activity that will probably generate vibrations in the E. M. Facility. I will try to determine if there will be likely quiet periods.
Chuck
This is what is on tap for the Vail electrical project for the next two weeks.

11/20-11/24:
Contractor will be completing the shifting off the road and sidewalk and the temporary paving on 11/20.
For the remainder of the week, they will be stripping the top soil next to the Remsen building. They are not working on 11/23 or 11/24. The noise/vibrations should be minimal.

11/27-12/1:
The contractor will be bringing in equipment to begin "driving" the sheet pilings needed to complete their excavation. They plan on beginning the actual driving on Tuesday. This process will create a fair amount of noise/vibrations. The entire process is scheduled to be completed by 12/8.

This is all that I have for now. I will keep you posted if anything changes.

Jon Rader
DMS Facilities

July 24, 2006.

For the next three to four weeks there will be construction activity outside the lab that will likely cause vibration problems. Starting Thursday or Friday of this week a trench and vault will be dug from the new dorms down the west side of Remsen. This is for a new power line. That work will take three or four weeks to complete. At the same time, there will be construction on a new road between Gilman and the new dorms. When the vibrating rollers start up, microscopy will be difficult. Please plan accordingly. I will try to keep you updated on the construction.
CPD

July 17, 2006.

Louisa is on vacation through July 21. Kate will be out on vacation from July 24 to August 1. Chuck will be at the Botanical Society of America meeting the week of July 31, and will be on vacation from August 7 through 16.
CPD

July 11, 2006.

A page listing fees has been added to the web site.
We will be having service on the Tecnai on July 12 to correct the persistent problems with the CompuStage. Several electronics boards will be replaced in an effort to fix things.
CPD

December 22, 2005.

Some general notes.
Next week we will have a service call on the Tecnai to correct a persistent stage problem.
January 20-24 we will have service on the SEM to install a new FEG tip, and to tune up the microscope. It will be ready for use at 1 pm Jan. 24, if no problems are encountered.
Thirsday and Friday nights on Jan 5 and 6 we will have floors cleaned in the microscope rooms. Please do not planon using the Facility between 10 pm and 6 am those days.
CPD

September 30, 2005.

After sending out my last message with the signup rules, I realized that I had not told you all of the rules. Not wanting to be accused of using Chicago rules (first rule is don't tell them all of the rules of the game), here is a fuller rule set. I have also modified the rules slightly after discussion with some groups.

Signups more than 7 days in advance will generate a warning, more than 14 days in advance will be rejected.
You may have up to 2 signups in advance as primary user, up to 4 in advance as alternate user. If you signup as an alternate, you should email the primary user so the will know there is an alternate. If you are the primary user and cancel (see below), you should contact the alternate (the program should prompt you to do so).
Cancellations less than 24 hours in advance may result in charges for 1/2 the normal charge, unless there is a good reason. Contact me and plead your case. I will be judge of what is a good reason.

In addition to these rules, on the SEM we need to keep the noon to 1 pm slot open for users who need to only check specimens, so do not schedule over this time (10-2 pm, for example) unless you check with me and convince me it is essential.

If the rules cause problems, let me know. They are a work in progress, and are intended to encourage efficient use of the instruments for many users. I am trying to get the scheduling program modified so each instrument can have its own set of rules, but for now all rules apply to all instruments.
CPD

September 26, 2005.

The online scheduling system will switch to hard enforcement of scheduling rules starting today. All violations will result in rejection of the attempted scheduling. The rules are:
Minimum time = 0.5 hours;
Maximum time = 4 hours - more than this will generate a warning.
Reservations more than 16 hours will be rejected.
Reservations more than 8 days in advance will generate a warning.
Reservations more than 15 days in advance will be rejected.
Accounts will expire after 120 days of inactivity. Contact the Director to re-activate accounts.
If the rules need to be relaxed for you, contact the Director and plead your case.
CPD

August 10, 2005.

I will be away on vacation from 8-11 until 8-29. If you need microscopy help, please contact either Louisa Howard or Kate Connolly.
CPD

July 19, 2005.

A power outage last night shut down the REMF server and several microscopes. The XL-30 stayed on, due to its backup power supplies. The Tecnai and the 1010 both shut down. The Tecnai has fully recovered, although we are waiting for the vacuum to improve with the holder in place. The 1010 is still pumping (its is almost there after 4 hours), and will be back up before the end of the day.

CPD

May 18, 2005.

This from the dorm project.

Dear McLaughlin Cluster Neighbors,
I am pleased to inform you that piledriving operations have finished, as of 11:30AM this morning. I can assure you that everyone is relieved to have that portion of the construction completed.
Sincerely,
Andrew Ager

CPD

April 28, 2005.

You may have noticed that the url for the EM Facility web site has changed. It is now http://www.dartmouth.edu/~emlab. Gone is ~emczar. Please change bookmarks . A redirect will be active for a short while.
CPD

April 26, 2005.

The online resource scheduling system is now active. Authorized users may now sign up for instruments using a web interface. Other users may check the schedule and request a session with Facility staff. Go to the Online Resource Scheduler now.
CPD


March 16, 2005.

There is a construction update web site.

I suggest you check it, rather than having me repeat what they post. Regardless of the scheduled activity, plan on intermittent vibrations, sometimes severe. AFM does not appear to be affected. All other lab activities will suffer, some worse thanothers.
CPD

Uploaded Image: aperture-3.jpg

Instrument Schedules